translate this to a mathematical equation and solve. please show all of your work
A football team gained 28 yards on one play, lost 33 yards on another, and gained 33 yards on the last play of the first half, they had already gained 207 yards during the half. what was the total yardage gain for the first half?

Answer by Earlsdon(6294)

Remember, a gain is + and a loss is -.
Gained 28 yards = +28
Lost 33 yards = -33
Gained 33 yards = +33
Gained 207 yards = +207 yards.
Add 'em all up.
28-33+33+207 = 28+207 = 235 yards total yardage gained.